# Beyond boundaries: Gary Lorden’s groundbreaking contributions to sequential analysis

### v1

@@ 3 Sequential Changepoint Detection: Lorden’s Minimax Change Detection Theory
where ${\mathsf{ESEDD}}_{\theta}(T)=\sup_{0\leqslant\nu<\infty}\operatornamewithlimits{ess\,sup}{\mathsf{E}}_{\nu,\theta}[(T-\nu)^{+}|{\mathscr{F}}_{\nu}]$ and ${\mathsf{E}}_{\nu,\theta}$ is the expectation under probability measure ${\mathsf{P}}_{\nu,\theta}$ when the change occurs at $\nu$ with the post-change density $f_{\theta}$.
@@
where $\theta_{1}$ may be either a fixed value if the alternative hypothesis is $\theta\leqslant-\theta_{1}$ or $\theta\geqslant\theta_{1}$ or $\theta_{1}(\gamma)\to 0$ as $\gamma\to\infty$ if the hypothesis is $\theta\neq 0$. Lorden shows that
@@
The inequality (78) is usually overly pessimistic. A much better result gives the approximation
${\mathsf{P}}_{0}(\tau(h)<\infty)\approx\sqrt{h(\gamma)}\exp\left\{-h(\gamma)\right\}C,$
which follows from (56). However, the latter one does not guarantee the inequality ${\mathsf{P}}_{0}(\tau(h)<\infty)\leqslant\gamma^{-1}$, and therefore, the inequality ${\mathsf{E}}_{\infty}[\tau^{*}(h)]\geqslant\gamma$.
@@
1. 15 years later Moustakides (1986) improved Lorden’s asymptotic theory by proving, using optimal stopping theory, that for any ARL to false alarm $\gamma\geqslant 1$ the CUSUM procedure is strictly optimal if threshold $a=a(\gamma)$ is selected so that $\operatorname{\mathsf{ARLFA}}(T_{a})=\gamma$.
2. Shiryaev (1996) showed that the CUSUM procedure is strictly optimal in the continuous-time case for detecting the change in the mean of the Wiener process for Lorden’s minimax criterion.
3. Pollak (1985) introduced a different minimax criterion of minimizing the supremum expected detection delay $\sup_{\nu\geqslant 0}{\mathsf{E}}[T-\nu|T>\nu]$ and a modification of the conventional Shiryaev–Roberts (SR) procedure (the SRP procedure) that starts from the random point distributed with the quasi-stationary distribution of the SR statistic and proved that it is third-order asymptotically minimax minimizing the $\sup_{\nu\geqslant 0}{\mathsf{E}}[T-\nu|T>\nu]$ to within $o(1)$ as $\gamma\to\infty$ in class ${\mathbb{C}}(\gamma)$.
4. Tartakovsky, Pollak, and Polunchenko (2012) proved that the specially designed SR-$r$ procedure that starts from a fixed point $r=r(\gamma)$ is third-order asymptotically optimal with respect to Pollak’s measure $\sup_{\nu\geqslant 0}{\mathsf{E}}[T-\nu|T>\nu]$ in the class ${\mathbb{C}}(\gamma)$ as $\gamma\to\infty$.
5. Polunchenko and Tartakovsky (2010) proved that the specially designed SR-$r$ procedure that starts from a fixed point $r=r(\gamma)$ is strictly optimal with respect to Pollak’s measure $\sup_{\nu\geqslant 0}{\mathsf{E}}[T-\nu|T>\nu)]$ in the class ${\mathbb{C}}(\gamma)$ for a particular model.
@@
7. Moustakides, Polunchenko, and Tartakovsky (2009) performed a detailed comparison of CUSUM and SR procedures which shows that the CUSUM is better in terms of the conditional expected detection delay ${\mathsf{E}}_{\nu}[T-\nu|\mid T>\nu]$ for relatively small values of the change point $\nu$ but less effective than the SR procedure for relatively large $\nu$.
8. Lorden and Pollak (2005, 2008) proposed adaptive SR and CUSUM procedures that exploit one-step delayed estimators of unknown parameters $\theta$ (an estimate $\hat{\theta}_{n-1}$ is used after observing the sample of size $n$) as in the Robbins–Siegmund one-sided adaptive SPRT (see Robbins and Siegmund (1972, 1974)) and compared their performance with that of the mixture-based SR procedure.
4 Consulting, Hollywood, and Major League Baseball
With his creativity and generosity of ideas, it is not surprising that Lorden was in demand as an academic collaborator and consultant, and much of his early work on changepoint detection arose from collaborations with researchers at the Jet Propulsion Laboratory (JPL) in Pasadena, California, which is managed for NASA by Lorden’s home institution of Caltech.
But Lorden was also an uncommonly effective, engaging, and entertaining communicator and this caused many outside academics to seek him out as a consultant too. Lorden would routinely be interviewed by reporters and appear on the evening news, explaining things like the odds of winning the latest lottery jackpot. Lorden also “moonlighted” as an expert witness in court cases involving statistics and mathematics.
Eventually Hollywood came calling, and in 2005 Lorden was asked to be the math consultant for a (then) new CBS television show called NUMB3RS. In a bit of art imitating life, the show was about a Caltech professor who used math to help the FBI solve crimes, and Lorden worked with the show’s writers to accurately incorporate mathematical topics into the storylines. An aspect of the relationship that Lorden particulalry relished, that he related to us, was that the show’s creators initially considered setting the show at MIT (Caltech’s rival), but decided to change the venue after learning more about Caltech and Lorden. The show would go on to be a hit, running for 118 episodes over 6 seasons. If a viewer of the show knew something of Lorden’s work they could see some of his favorite topics (and many topics in this survey article) woven into the episodes’ plot lines including changepoint detection, hypothesis testing, Bayesian methods, gambling math, cryptography, and sports statistics among others. With Keith Devlin, Lorden wrote a popular general audience book (Devlin and Lorden 2007) on the mathematical topics appearing in the show. For example, Devlin and Lorden (2007, Chapter 4) explain the basic concept of changepoint detection:
and goes on to discuss the importance of this method for quick response to potential bioterrorist attacks and for designing efficient algorithms to pinpoint various kinds of criminal activity, such as to detect an increase in rates of crimes in certain geographical areas and to track changes in financial transactions that could be associated with criminal activity.

### v2

@@ 3 Sequential Changepoint Detection: Lorden’s Minimax Change Detection Theory
where ${\mathsf{ESEDD}}_{\theta}(T)=\sup_{0\leqslant\nu<\infty}\operatornamewithlimits{ess\,sup}{\mathsf{E}}_{\nu,\theta}[(T-\nu)^{+}|{\mathscr{F}}_{\nu}]$ and ${\mathsf{E}}_{\nu,\theta}$ is the expectation under ${\mathsf{P}}_{\nu,\theta}$ when the change occurs at $\nu$ with the post-change density $f_{\theta}$.
@@
where $\theta_{1}$ may be either a fixed value if the alternative hypothesis is $\theta\leqslant-\theta_{1}$ or $\theta\geqslant\theta_{1}$ or $\theta_{1}(\gamma)\to 0$ as $\gamma\to\infty$ if the hypothesis is $\theta\neq 0$. Lorden demonstrates that
@@
The inequality (66) is usually overly pessimistic. A much better result gives the approximation ${\mathsf{P}}_{0}(\tau(h)<\infty)\approx\sqrt{h(\gamma)}\exp\left\{-h(\gamma)\right\}C,$ which follows from (44). However, the latter one does not guarantee the inequality ${\mathsf{P}}_{0}(\tau(h)<\infty)\leqslant\gamma^{-1}$, and therefore, the inequality ${\mathsf{E}}_{\infty}[\tau^{*}(h)]\geqslant\gamma$.
Later, Lorden and Pollak (2005, 2008) proposed adaptive Shiryaev-Roberts and CUSUM procedures that utilize one-step delayed estimators of unknown post-change parameters $\theta$. In these procedures, an estimate $\hat{\theta}_{n-1}(X_{1},\dots,X_{n-1})$ is used after observing the sample of size $n$, similar to the Robbins–Siegmund one-sided adaptive SPRT; see Robbins and Siegmund (1972, 1974). They compared the performance of these adaptive procedures with that of the mixture-based Shiryaev-Roberts procedure. Notably, these adaptive procedures are computationally simpler than the GLR CUSUM procedure.
@@
1. Fifteen years later, Moustakides (1986) advanced Lorden’s asymptotic theory by demonstrating, using optimal stopping theory, that the CUSUM procedure is strictly optimal for any ARL to false alarm $\gamma\geqslant 1$ if the threshold $a=a(\gamma)$ is chosen such that $\operatorname{\mathsf{ARLFA}}(T_{a})=\gamma$.
2. Shiryaev (1996) showed that the CUSUM procedure is strictly optimal in the continuous-time scenario for detecting the change in the mean of the Wiener process according to Lorden’s minimax criterion.
3. Pollak (1985) introduced a distinct minimax criterion aimed at minimizing the supremum expected detection delay $\sup_{\nu\geqslant 0}{\mathsf{E}}[T-\nu|T>\nu]$. Additionally, Pollak proposed a modification of the conventional Shiryaev–Roberts (SR) procedure known as the SRP procedure, which initiates from a randomly distributed point following the quasi-stationary distribution of the SR statistic. He proved that this procedure is third-order asymptotically minimax, minimizing $\sup_{\nu\geqslant 0}{\mathsf{E}}[T-\nu|T>\nu]$ to within $o(1)$ as $\gamma\to\infty$ within the class ${\mathbb{C}}(\gamma)$.
4. Tartakovsky, Pollak, and Polunchenko (2012) proved that the specially designed SR-$r$ procedure that starts from a fixed point $r=r(\gamma)$ is third-order asymptotically optimal with respect to Pollak’s measure $\sup_{\nu\geqslant 0}{\mathsf{E}}[T-\nu|T>\nu]$ within the class ${\mathbb{C}}(\gamma)$ as $\gamma\to\infty$.
5. Polunchenko and Tartakovsky (2010) demonstrated that the specially designed SR-$r$ procedure, which commences from a predetermined point $r=r(\gamma)$, is strictly optimal with respect to Pollak’s measure $\sup_{\nu\geqslant 0}{\mathsf{E}}[T-\nu|T>\nu]$ within the class ${\mathbb{C}}(\gamma)$ for a specific model.
@@
7. Moustakides, Polunchenko, and Tartakovsky (2009) conducted a thorough comparison of CUSUM and SR procedures, demonstrating that CUSUM outperforms SR in terms of the conditional expected detection delay ${\mathsf{E}}_{\nu}[T-\nu|T>\nu]$ for relatively small values of the change point $\nu$. However, SR proves to be more effective than CUSUM for relatively large $\nu$.
@@ Acknowledgements
We express our gratitude to the reviewers, the associate editor, and the editor for their valuable comments, which greatly enhanced the quality of the paper. We also thank Caltech for the use of the photos in Figure 1.